MIXED QUANTUM/CLASSICAL APPROACH TO PHOTODISSOCIATION OF H_{2}O AND Ar − H_{2}O

\underline{FENG CHEN} and ANNE B. MCCOY, Department of Chemistry, The Ohio State University, Columbus, OH 43210.

Recently, experimental studies of the vibrationally mediated photodissociation of H_{2}O in Ar − H_{2}O van der Waal complexes have been reported. In these studies, the OH rotational distributions were measured and compared with the ones that resulted from the photodissociation of H_{2}O at the same photolysis wavelengths.^{a} Much warmer OH rotational distributions were obtained from photolysis of the Ar − H_{2}O clusters. In order to understand the physical origins of these results we have simulated the photodissociation of H_{2}O and Ar − H_{2}O using a multiple configuration mixed quantum/classical approach. By comparing the rotational distribution of the OH products that are obtained from photodissociation of the complexed and uncomplexed water molecules, we find that the Ar atom has a strong effect on the rotational distribution of the OH fragment in intracluster collisions. The simulations provide insights to the dynamics of photodissociation in clusters.

aVotava, O.; Plusquellic, D. F.; Myers, T. L.; Nesbitt, D. J. J. Chem. Phys. \underline{112}, 7449 (2002).